OpenStudy (nainai100):

What multiply to 26 and add up to 7 ?

myininaya (myininaya):

xy=26 x+y=7 if xy=26 then y=26/x plugging this into our second equation we have x+26/x=7 multiply both sides by x x^2+26=7x x^2-7x+26=0 use quadratic equation to find x

jimthompson5910 (jim_thompson5910):

Make a table of all the factors (both positive and negative) 1 26 2 13 -1 -26 -2 -13 Now add up each pair 1+26 = 27 2+13 = 15 -1+(-26) = -27 -2+(-13) = -15 So there are no two rational numbers that multiply to 26 and add to 7

myininaya (myininaya):

\[x=\frac{7 \pm \sqrt{(-7)^2-4(1)(26)}}{2(1)}=\frac{7 \pm \sqrt{49-104}}{2}\] make that imaginary

OpenStudy (ybarrap):

a*b=26 a+b=7 b=26/a a + 26/a = 7 a^2 + 26 = 7*a a^2 - 7a + 26 = 0 7^2 - 4*26 = 49 - 104 < 0 so solutions will be complex. There are no real solutions.

myininaya (myininaya):

we have to look at all complex numbers to answer this one

myininaya (myininaya):

\[x=\frac{7 \pm \sqrt{-55}}{2}=\frac{7 \pm i \sqrt{55}}{2}\] y=26/x \[=\frac{26(2)}{7 \pm i \sqrt{55}}=\frac{52}{7 \pm i \sqrt{55}}\]

OpenStudy (nainai100):

never mind i messed that wasn't even a problem i looked at 2 different problems and put it as 1 sorry lol

OpenStudy (nainai100):

up*